# A Scoping Review of Synthetic Data Generation for Biomedical Research and Applications

### v1

@@ Discussion
This scoping review has several limitations that should be acknowledged to appropriately interpret our findings and analysis. First, the literature search was limited to specific databases and publication venues, potentially missing relevant studies published in less prominent journals or literature sources. Second, our review primarily included studies written in English, potentially overlooking important findings published in other languages. Additionally, the rapid pace of research in LLM-driven synthetic data generation means that emerging studies and recent methodological advances might not be captured fully in this review. Lastly, given the inherent heterogeneity in study designs, evaluation metrics, and application domains across the reviewed literature, direct comparison and generalization of findings are challenging, underscoring the need for more standardized reporting and evaluation frameworks in future research.
Methods
Data Source and Search.
We conduct searches between April and May 2025 following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines [73] and the PRISMA Extension for Scoping Reviews (PRISMA-ScR) [74] (Figure 2). Recognizing the rapid developments of Large Language Models (LLMs) in the biomedical field, our review encompassed peer-reviewed articles between January 1, 2020, to April 5, 2025 (inclusive) from multiple database sources, including ACM Digital Library, PubMed, Web of Science, and Google Scholar. Supplemental articles were gathered by reviewing article bibliographies and by soliciting suggestions from XH, HW, IH, and ZH. Our search strategy utilized a logical combination of keywords: “Language Model” AND “synthetic data” AND “medical OR health OR clinical”. We primarily limited the search to titles and abstracts, expanding to full texts where this function was unavailable. For Google Scholar, we conducted a keyword search, sorted by relevance, and selected the first 100 studies. The search strategies for each database were initially formulated in the early stages of the study and refined through team discussions and preliminary analysis of the results.
Figure 2: PRISMA-ScR flow diagram.
Study Selection.
We followed inclusion and exclusion criteria to narrow the candidate articles. Titles and abstracts were initially screened against predefined eligibility criteria by at least two independent reviewers, excluding records that clearly did not meet inclusion criteria and retaining ambiguous ones for full-text screening. Our inclusion criteria include: 1) an article conducts biomedical research; 2) an article uses Large Language Models (LLMs) for synthetic generation of biomedical data; and 3) an article has undergone peer review and been published in a journal or in conference proceedings within the range of 01/01/2020 and 04/05/2025. Our exclusion criteria include: 1) an article is a survey, literature review, news article, editorial, letter, opinion, study protocol, and comment; 2) an article is not in English; 3) articles do not study biomedical large language models; and 4) an article is not peer-reviewed and published in a journal or conference proceedings. Three undergraduate students from the XH lab assisted with review of candidate articles with the HS and XH, who divided the candidate articles into relatively equal numbers and assigned them to reviewers. Reviewers examined titles, abstracts, full-text, and other article metadata, recommending inclusion or exclusion based on the defined criteria. Each article was reviewed by at least two individuals, and disagreements were resolved after further review by XH and HW.
Data Extraction.
Selected articles were mainly processed by two individuals (HR and XH), who are experts in natural language processing (NLP) and biomedical informatics. They independently reviewed the full text of each selected article and extracted key metadata. The extracted data focused on three main aspects: 1) Synthetic data types and applications, including data characteristics , such as language and domain; 2) Synthetic data generation methods, presenting generation pipelines (models used to generate synthetic data) and their data modalities (e.g., text, image, audio); 3) Quality assessment and evaluation metrics summarizing medical applications, evaluation approaches, and evaluation settings. Discrepancies in data extraction were resolved through discussion between HR and XH, with further arbitration by XH, HW, and IH when necessary. All extraction processes were documented for replication purposes.
Acknowledgments
@@
WR, WL, and WH were partially supported by the National Science Foundation under Award IIS-2245920 and IIS-2440381, and the work was partially supported by National Cancer Institute Award R01CA258193. ZH was supported by the Agency for Healthcare Research and Quality Award R21HS029969, the National Library of Medicine Award R21LM013911, the National Institute of Mental Health Award R21MH137736, and the National Institute on Aging Award R01AG064529. The authors would like to thank three undergraduate students, Sampoorna Poria, Sashank Reddy Karri, and Precious T Jones, for assisting paper collection and filtering at the initial stage.
Author contributions statement
@@
Additional information
Competing interests
The authors declare no competing interests.
Data availability
Not applicable.

### v2

@@ 4 Discussion
This scoping review should be interpreted in light of several constraints. First, the search was restricted to four major databases (ACM Digital Library, PubMed, Web of Science, and Google Scholar) and to peer-reviewed venues; grey literature, non-indexed conference proceedings, and pre-prints may therefore be under-represented. Second, we limited inclusion to studies published in English between 1 January 2020 and 5 April 2025, so potentially relevant work in other languages or outside this time frame was not captured. Third, consistent with the PRISMA-ScR framework, we performed no critical appraisal of individual study quality; consequently, the methodological robustness of the included papers was not formally assessed. Additionally, the heterogeneity of study designs, synthetic-data tasks, evaluation metrics, and biomedical application domains precluded quantitative synthesis and limits the generalizability of pooled observations. Reproducibility is further limited by reliance on proprietary, non-deterministic APIs and incomplete reporting of prompts/decoding settings; hallucination audits were inconsistently performed across included studies. Lastly, given the rapid pace of innovation in LLM-based synthetic data generation, newly published methods appearing after our search window may not be reflected here, underscoring the need for periodic updates of this evidence map.
5 Conclusion
Our scoping review examines the current trends and challenges of utilizing synthetic data generated by language models in biomedical fields with a focus on utility and quality aspects. The results indicate that to fully utilize the LLM-generated synthetic data, the field may address several critical issues, including standardized evaluation frameworks, increased transparency, human-AI collaboration, and accessible models and data. Although our review was limited by the rapid pace of advancements and a primary focus on English-language publications, these challenges present important opportunities for collaboration and innovation. By promoting reproducibility, rigorous evaluation, and expert feedback for the synthetic generation approach, our study shows the integration of high-quality synthetic data can benefit biomedical research and clinical practice.
Statements and Declarations
Funding
@@ 5 Conclusion
HR, WL, and XH were partially supported by the National Science Foundation under Award IIS-2245920 and IIS-2440381, and IH was partially supported by National Cancer Institute Award R01CA258193. HW was supported by NIH Office of Strategic Coordination (Common Fund) R03OD038389. ZH was supported by the Agency for Healthcare Research and Quality Award R21HS029969, the National Library of Medicine Award R21LM013911, the National Institute of Mental Health Award R21MH137736, and the National Institute on Aging Award R01AG064529.
Conflict of interest
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Data availability
Data are available upon request to the corresponding author.
Materials availability
Not applicable.
Code availability
Not applicable.
Author contribution
@@
Appendix A Database-specific search strategies
Database Search string
PubMed (“large language model*”[Title/Abstract] OR “LLM*”[Title/Abstract]) AND (synthetic*[Title/Abstract]) AND (data*[Title/Abstract]) AND (biomedical*[Title/Abstract] OR medical*[Title/Abstract] OR clinical*[Title/Abstract] OR health*[Title/Abstract])
ACM Digital Library Abstract:(“large language model*” OR “LLM*”) AND Abstract:synthetic* AND Abstract:data* AND Abstract:(biomedical* OR medical* OR clinical* OR health*)
Web of Science TS=(“large language model*” OR “LLM*”) AND TS=(synthetic*) AND TS=(data*) AND TS=(biomedical* OR medical* OR clinical* OR health*)
Google Scholar (“large language model” OR “large language models” OR LLM) AND synthetic AND data AND (biomedical OR medical OR clinical OR health)
Table 2: Search strategies for each database included in this review.